Simmons Marketing (SM) has the capacity to produce 10,000 fax machines per year. SM currently produces and sells 7,000 units per year. The fax machines normally sell for $100 each. Modem Products has offered to buy 2,000 fax machines from SM for $60 each. Unit-level costs associated with manufacturing each fax machine are $15 for direct labor and $40 for direct materials. Product-level and facility-sustaining costs are $50,000 and $65,000, respectively.

1. What is SM's current profit (net income)?

a. $115,000

b. $120,000

c. $200,000

d. $315,000

2. How much would profit increase (decrease) if SM accepted this special order?

a. $10,000

b. $112,000

c. ($10,000)

d. ($112,000)
Answer #1

1. Sales revenue = 100*7000 = 700,000

Variable cost = (15+40)*7000 = 385,000

Fixed cost = 50,000+65,000 = 115,000

Net income = 700,000-385,000-115,000

= 200,000

Option C is the answer

.

2. Relevant costs = (15+40)*2000 = 110,00

Special order revenue = 2000*60 = 120,000

Increase in profit = 120,000 - 110,000 = 10,000

Option A is the answer